sam has 50 nickels and quarters worth $8.90
how many quarters does he have?
please show how you got your answer

Answer by stanbon(75887)

sam has 50 nickels and quarters worth $8.90
how many quarters does he have?
----
Quantity Equation: n + q = 50
Value Equation::: 5n+25q = 890
--------------------------------
Multiply thru the Quantity equation by 5:
5n + 5q = 5*50
5n + 25q= 890
-------
Subtract and solve for "q":
20q = 640
q = 32 (# of quarters)
----
solve for "n":
n + q = 50
n + 32 = 50
n = 18 (# of nickels)
=========================
Cheers,
Stan H.
=========================